OpenStudy (marinadanyel):

What are the vertical asymptotes of the function f(x) = the quantity of 3x plus 9, all over x squared plus 4x minus 12 ?

OpenStudy (marinadanyel):

\[\frac{ 3x+9 }{ x^2+4x-12 }\]

OpenStudy (marinadanyel):

Answer choices: x = −6 and x = −2 x = −6 and x = 2 x = 1 and x = −2 x = 1 and x = 2

OpenStudy (marinadanyel):

Do I factor here?

OpenStudy (photon336):

\[\frac{ 3x+9 }{ x^{2}+4x-12 }\] try factoring canceling and setting the denominator to zero. \[\frac{ 3(x+3) }{ (x+6)(x+2) }\]

OpenStudy (marinadanyel):

Okay, from what I see, nothing can cancel here?

OpenStudy (photon336):

next let's look at the denominator and split it up

OpenStudy (marinadanyel):

so the vertical asymptotes are x=-6 & x=-2 ???

OpenStudy (photon336):

can you solve these? (x+6) = 0 and (x+2) = 0

OpenStudy (marinadanyel):

Yeah, so I solve for x, right?

OpenStudy (photon336):

yep

OpenStudy (marinadanyel):

Thanks!

OpenStudy (marinadanyel):

What about this one?

OpenStudy (photon336):

hey let me show you the graph and how it looks

OpenStudy (marinadanyel):

What is the oblique asymptote of the function f(x) = the quantity of x squared plus 7 x plus 11, all over x plus 5 ?

OpenStudy (marinadanyel):

What's the difference between vertical asymptotes & oblique asymptotes?

OpenStudy (marinadanyel):

How would I solve?

OpenStudy (photon336):

http://prntscr.com/bjcexh

OpenStudy (marinadanyel):

\[\frac{ x^2+7x+11 }{ x+5 }\]

OpenStudy (photon336):

hey @marinadanyel can we close this question and open up another one?

OpenStudy (marinadanyel):

Yes, but now I am curious about that graph.

OpenStudy (photon336):

http://prntscr.com/bjcexh

OpenStudy (photon336):

take a look

OpenStudy (marinadanyel):

I saw, but how does that work?

OpenStudy (photon336):

how does what work?

OpenStudy (marinadanyel):

Can they touch if they are vertical asymptotes?

OpenStudy (photon336):

no

OpenStudy (photon336):

to my knowledge the graph will approach that value but never touch it.

OpenStudy (marinadanyel):

Oh, okay I see now.

OpenStudy (marinadanyel):

I'll open a new question.

OpenStudy (radar):

The denominator factors need to be verified, reviewed etc.

OpenStudy (marinadanyel):

What?

OpenStudy (radar):

(x + 6)(x + 2) does not equal x^2 + 4x - 12

OpenStudy (radar):

Close....but no cigar!

OpenStudy (marinadanyel):

Oh, you're right.

OpenStudy (marinadanyel):

Nice! Thanks so much, it should be \[\frac{ 3(x+3) }{ (x-2)(x+6) }\]

OpenStudy (marinadanyel):

Correct? Giving me x = 2 & x = -6 instead.

OpenStudy (radar):

Hope that checks out, that is what I thought it should be.

OpenStudy (robtobey2):

Refer to the attachment